A particle execute simple harmonic motion about the point x = 0. At t= 0, it has

displaced y = 0.37 cm, and zero velocity. If the frequency of motion is 0.25/s. Calculate

the period, amplitude, the maximum speed and the maximum acceleration.


1
Expert's answer
2021-02-10T10:05:41-0500

The equation of SHM in general form is:


"y(t) = A\cos(\omega t)"

where "A = 0.37cm" is the amplitude, "\omega = 2\pi f = 2\pi\cdot0.25rad/s = 0.5\pi \space rad/s" is the angular frequncy. The period is:


"\omega T = 2\pi\\
T = \dfrac{2\pi}{\omega} = 4s"

The speed is (and it is 0 at t=0):


"\dot{y}(t) = -A\omega\sin(\omega t)"

And the maximum speed is:


"\dot{y}_{max} = A\omega = 0.175\pi \space cm/s"

The acceleration is:


"\ddot{y}(t) = -A\omega^2\cos(\omega t)"

And the maximum acceleration is:


"\ddot{y}_{max} = A\omega^2 = 0.37\cdot (0.5\pi)^2 =0.0925\pi^2 \space cm/s^2"

Answer. Period: "4s" , amplitude: "0.37cm", maximum speed "0.175\pi \space cm/s", maximum acceleration: "0.0925\pi^2 \space cm/s^2"